# How Haag-tied is QFT, really?

### 18 Dec 2023 v2

@@ 5.1 Three distinct lines of debate
The most salient of disagreements concerned the status of the interaction picture. On the one hand, Earman and D. Fraser, Klaczynski, and Maiezza and Vasquez each suggest its existence is undermined by Haag’s theorem. On the other hand, Duncan and Miller both suggest it exists and its use is justified. The difference here is, of course, that Duncan and Miller speak of the interaction picture at intermediate stages in a perturbative calculation, where regularization ensures finitely-many degrees of freedom. In those contexts, the interaction picture is perfectly well defined. Our framework was not necessary for this observation, nor was an appeal to extra-Haagian outlooks.
Yet there is more to say here, as revealed by the framework. First, the two groups differ substantially on what QFT should look like and, therefore, what counts as a sure foundation. Duncan and Miller are focused first and foremost on the reliability of the predictions generated by “the algorithm of quantum field theory” [Miller, 2021, 1135]. While Miller clearly believes a non-perturbative structure is desirable [Miller, 2018, 818], he nevertheless argues that the perturbative nature of our data does not undermine a realist interpretation of the theory [Miller, 2021, 1135]. Thus, if the ultimate goal is to characterize a theory about which we can be realists, it would seem canonical QFT has a sufficiently sure foundation. Earman and D. Fraser, Klaczynski, and Maiezza and Vasquez do not seem similarly satisfied with canonical QFT’s present perturbative structure.
Second, members of the former group are not entirely alike in their accounts of the consequences of Haag’s theorem. At issue is how deeply the authors trace Haag’s theorem’s consequence for the existence of an intertwiner. For Earman and D. Fraser, the focus is on the IP as an intertwiner between the free field and the unrenormalized interacting field. For Klaczynski and Maiezza and Vasquez, it is on the IP as an intertwiner between the free field and the renormalized interacting field. The differing focuses lead to different proposed strategies for renovating QFT.
@@
The reader might think Kastner fits in with Bain. This isn’t correct, however. Kastner not only recommends a different interpretation of particle physics from Halvorson and Clifton but, as her proposed renovations reveal, suggests QFT must be replaced altogether. That is, rather than derive an interpretation from a presupposed formalism, she instead derives a formalism from a presupposed interpretation. Among other things, this interpretation is committed to a particle conception of particle physics. This commitment ultimately derives from a belief that relativistic quantum mechanics should behave more non-locally than QFT does, as, she thinks, non-relativistic quantum mechanics counsels. Moreover, there is at least one respect in which Kastner fits more easily with Halvorson and Clifton: despite its (supposed) empirical equivalence to DATs, QFT is to be replaced. This suggests two disanalogies with Bain. First, the comparatively unmoored approximations of canonical QFT are not enough—we should demand exact calculations, or at least approximations to exact solutions. Second, a fully satisfactory theory should wear its metaphysics on its face more plainly than does canonical QFT.
@@
Spelling out the renovations proposed also reveals commitments we might have missed. Focusing just on broad-strokes aims as a way to separate various responses to Haag’s theorem, Kastner and Seidewitz would seem far removed from the pristine interpreters—the former propose new theoretical forms for QFT, while the pristine interpreters restrict themselves to the formal variants. However, it is worth noting that these authors seem to share the same outlook on axiomatization. In particular, both seem at least weakly committed to claiming that QFT phenomena has a structure that can be captured in a closed-form as something like a finite set of axioms and that we can actually specify these axioms. Likewise, the pristine interpreters’ project is predicated on a belief that such a form exists and that we have found it. In contrast, Earman and D. Fraser appear committed only to the belief that axiomatization could be helpful for comprehending current QFT, or even for finding a more scrutable form. D. Fraser, at least, is quite clear that CQFT does not need to build a realistic model in order to have been useful—its utility lies rather in the fleshing-out of the relationship among the axioms. While Kastner and Seidewitz certainly could take such a view, the emphasis they place on the correctness of their chosen axioms suggests this is not their current conception.
@@
This is useful, too, for understanding the varied responses to the question of which assumption of Haag’s theorem ‘should’ be dropped. Kastner, Seidewitz, and the pristine interpreters each have a clear answer to this question. Things get murkier as we consider the others, however. Not only do none of them explicitly affirm some set of precise assumptions, they are not entirely clear about which assumptions are problematic and how. This comes in degrees, of course. Earman and D. Fraser are perhaps clearest, suggesting that the IP’s assumption of a global unitary intertwiner is problematic but that the local equivalence delivered by Haag-Ruelle appears to suffice. Klaczynski and Maiezza and Vasquez are a bit less clear. Insofar as they each hope for an explicit formalism, unlike canonical QFT, they appear sympathetic to axiomatic approaches. Yet Klaczynski only suggests that renormalization somehow replaces the unitarity of the intertwiner, and Maiezza and Vasquez are even less precise about what is problematic. In a sense, Duncan is even less clear, albeit for a principled reason. Obviously respectful of axiomatization’s utility, he nevertheless emphasizes again and again that the truly remarkable feature of QFT is the “scale separation” property of local field theory [Duncan, 2012, 570-1]. Building one’s theory around this fact, as the effective Lagrangian approach of canonical QFT has done, would seem to undermine specification of any single set of assumptions since, e.g., symmetry structure can vary. Duncan, it would seem, cannot be clearer about what is denied.
@@ 5.2 Lessons for making progress
The framework’s breakdown of responses into assessments, repairs, and renovations and consequent revealing of extra-Haagian outlooks has thus distinguished the foregoing three separate essential questions raised by Haag’s theorem. Are the machinations of the framework strictly necessary for pulling apart these distinctive debates? For the respondents we review, perhaps not: they are each quite explicit about their goals and, therefore, about how fundamental an evasion of Haag’s theorem is to their goal. But not all responses, especially those that engage only glancingly with Haag’s theorem, will be or can be so explicit. To be sure, alternative frameworks could be used for sorting out the various viewpoints on Haag’s theorem. We could, for instance, simply generate a list of different questions various researches ask, with Haag’s theorem playing some role in their answers. This strategy, however, is likely to miss out on the connections between what we have called assessments and the longer term work of maintenance and renovation. While we give no argument as to the superiority of our framework above all others, nevertheless, we hope that the reader has found our framework to be of positive epistemic utility in the breadth of issues it captures and in the dialectical organization it delivers.
We began with the question, how Haag-tied is QFT, really? That is, how big of a problem for QFT is the problem (or problems) identified by Haag’s theorem? Clearly, there is no straightforward answer to that question. How one thinks about whatever problems Haag’s theorem raises, how those problems ought to be solved, and what future major work is called for in this area, depends heavily on one’s extra-Haagian outlooks. The reason why the literature on Haag’s theorem initially presents as a dizzying conceptual labyrinth (recall section 3), prior to some framework or other regimenting the various responses to the theorem, is that the extra-Haagian outlooks are the driving forces in this dialectic. Thus, the discourse on Haag’s theorem isn’t so much about Haag’ theorem, as it is about what hopes and expectations each member of the debates has for QFT.
Bringing these extra-Haagian outlooks out of the background and into the foreground sheds light on deeper commitments about what goals are well-suited for making progress in physics, as well as convictions about which methods are most apt to succeed at achieving those goals. Interestingly, despite the many disagreements amongst authors discussed in this paper, one central agreement about hoped-for progress stands at the end of the day. Specifically for QFT, all agree that finding the unambiguous, non-perturbative structure underlying canonical QFT would be enthusiastically welcomed progress in this area of physics. In other words, it would be progress if we could identify an exact, non-perturbative solution.
In the meantime, however, accepting that such an unambiguous, non-perturbative underlying structure has not yet been developed, how significant of an impediment to progress on smaller-scale research objectives is Haag’s theorem? That is, how much of an impediment is Haag’s theorem to “expanding the rooms” of QFT? By now it will come as no surprise that the answer varies, and that it varies according to the “expansions” respondents have in mind (see Figure 1 for a summary). Since it presents no impediment to building a tower of effective field theories or producing asymptotic calculations, Haag’s theorem is not a fundamental impediment to Duncan and Miller, nor to Wallace or Bain. In one sense, it is also not problematic for Kastner and Seidewitz—after all, QFT incorrectly represents quantum (resp. relativistic) principles anyways. Of course, Haag’s theorem would be an impediment if we were wedded to the scaffolding around which we have built the room, i.e., the assumptions of Haag’s theorem. For Earman and D. Fraser it no longer presents a significant barrier to representing interactions, since Haag-Ruelle and CQFT offer conceptually coherent alternative approaches to scattering theory. Indeed, “[a]xiomatic and constructive field theorists have digested the moral [of Haag’s theorem] and moved on to other problems”[Earman and Fraser, 2006, 334].
Haag’s theorem seems to be a more significant impediment for other proposed expansions. For Klaczynski and Maiezza and Vasquez, the desired expansion is a non-perturbative, unambiguous mathematical structure supporting canonical QFT. Circumventing Haag’s theorem is fundamental to this expansion, and appeal to CQFT and Haag-Ruelle is not an option since neither can support canonical QFT’s calculations. It also presents a substantial problem for pristine interpreters and those who desire a particle interpretation of interactions rooted in the free-field Fock space notion of particle. It is less of a problem for those open to weaker or non-fundamental, emergent conceptions of particle, like Bain, Wallace, or Feintzeig.
One is tempted to ask: how fundamental Haag’s theorem is really? As we construe it here, fundamentality is relative to an imagined expansion. But since which expansion one imagines depends on what one expects of a theory, Haag’s theorem’s fundamentality depends ultimately on one’s extra-Haagian outlook. Thus, to ask how fundamental Haag’s theorem is really is to ask which extra-Haagian outlooks is most reasonable. And as we saw in response to ‘What is relativistic QFT?’, agreement on a goal doesn’t necessarily settle all methodological questions. In particular, answers to the following can still differ:
@@
    What role does (should) foundational work play in progress in physics?
@@
    How does (should) foundational work coordinate with non-foundational work?
@@
    And, what does (should) foundational work even look like?
These are not new questions: the so-called Wallace-D. Fraser debate concerns exactly these kinds of questions [Wallace, 2006, Wallace, 2011, Fraser, 2011, Fraser, 2009b] (see also [Koberinski, 2016, Koberinski, 2022] for reception of the debate). Nevertheless, their implication here again highlights the utility of more careful assessments of answers than have been provided. For example, what does the relevant history suggest we can reasonably expect from axiomatic-adjacent approaches, and how does that compare to what is, in fact, being expected of them? Some work has been done here. For example, [Koberinski, 2023, 14] has argued that the axiomatic-adjacent approach of framework generalizations is reasonable insofar as it “expand[s] the domain of indirect tests able to find evidence of new physics.” In another vein, [Mitsch, 2022] argues that no-go theorems can anchor mathematical research programs in serious interpretive issues. [Freeborn et al., 2023] argue, moreover, that no-go theorems can serve as helpful guides to building adequate models. Further work in this area is needed.
The importance of Haag’s theorem is, in fact, an excellent example of the importance of these foundational questions. Recall that many folks agree that finding an unambiguous structure underlying canonical QFT’s perturbative calculations would be significant progress. Insofar as finding this structure would necessitate identifying the analytic structure to which the perturbative calculations are approximations, the phenomena of Haag’s theorem are directly implicated. Yet even with Haag’s theorem being so ‘fundamental’ in this sense, it is unclear how central these phenomena should be in our efforts to find this unambiguous structure. Indeed, perhaps its identification will come by completely ignoring Haag’s theorem and other foundational issues in favor of performing more high-energy experiments. The questions above, then, suggest we should concern ourselves with what method(s) we can reasonably expect to get us closer to this structure.
“Expanding the rooms” of QFT may not, in the end, require answers to these questions. That is, we may identify a generally acceptable expansion of QFT without any further reflection or clarity on the role of foundational work. Indeed, such an expansion could even determine our path out of the Haagian labyrinth. But this does not mean answering these questions would be useless. At worst, a more robust understanding of the methodological tradeoffs could provide heuristic guidance to individual researchers. Better still, it could focus effort on methods more likely to bear fruit, and it could help foster the clarity and humility needed to change course when one’s preferred method is no longer making progress. At best, it could identify the single best method for a given goal. In any case, it is worth gathering more evidence on the past success of foundational views and their prospects for success today.
Different extra-Haagian outlooks shape different long term goals for the development and interpretation of QFT, which in turn lead to different types of primary disciplinary long term renovation work. And yet, if work on QFT in physics, mathematics, and philosophy continues to aspire to be ultimately about the same part of the edifice of science, then scholarly discourse on results such as Haag’s theorem needs to have clear and efficient cross-disciplinary communication. How might we all—mathematicians, physicists, and philosophers—best direct our efforts in a well-coordinated pursuit of our various inter-connected goals? As summarized in figure 2, we can readily (in schematic, simplified form) sort proposals for potential methodologies to employ in service of these goals according to primary disciplinary homes. It is not, however, enough to advocate for one’s preferred methodology in the absence of a substantive and detailed argument in favor of that methodology’s likelihood for success. We might fruitfully draw here upon Ruetsche’s advocacy for non-pristine theory interpretation: just as context matters for explanatorily adequate theory interpretation, context matters for the reliability of inter-disciplinary methods. We therefore urge all those with a stake in the future of QFT to give serious thought to our three foundations-questions. These foundations-questions are policy generating questions, and good policy requires precisely-characterized methods with clear and persuasive evidence as to their sufficiency for solving the problem at hand. A high degree of specificity is necessary for making wise judgments about how to optimize the interdisciplinary community’s distribution of resources amongst these methodologies. QFT is not an epistemic island: there is a rich history of physics—and indeed of other sciences!—on which we can draw for evidence concerning the specific conditions in which a given methodology is likely (or not) to succeed in pursuit of goals that implicate the foundations of a given theory.

### 06 Apr 2024 v3

@@ 5.1 Three distinct lines of debate
The most salient of disagreements concerned the status of the interaction picture. On the one hand, Earman and D. Fraser, Klaczynski, and Maiezza and Vasquez each suggest its existence is undermined by Haag’s theorem. On the other hand, Duncan and Miller both suggest it exists and its use is justified. The difference here is, of course, that Duncan and Miller speak of the interaction picture at intermediate stages in a perturbative calculation, where regularization ensures finitely many degrees of freedom. In those contexts, the interaction picture is perfectly well defined. Our framework was not necessary for this observation, nor was an appeal to extra-Haagian outlooks.
Yet there is more to say here, as revealed by the framework. First, the two groups differ substantially on what QFT should look like and, therefore, what counts as a sure foundation. Duncan and Miller are focused first and foremost on the reliability of the predictions generated by “the algorithm of quantum field theory” [Miller, 2021, 1135]. While Miller clearly believes explicit identification of a non-perturbative structure is desirable [Miller, 2018, 818], he nevertheless argues that the perturbative nature of our data does not undermine a realist interpretation of the theory [Miller, 2021, 1135]. Thus, if the ultimate goal is to characterize a theory about which we can be realists, it would seem canonical QFT has a sufficiently sure foundation. Earman and D. Fraser, Klaczynski, and Maiezza and Vasquez do not seem similarly satisfied with canonical QFT’s present perturbative structure.
Second, members of the former group are not entirely alike in their accounts of the consequences of Haag’s theorem. At issue is how deeply the authors trace Haag’s theorem’s consequence for the existence of a unitary intertwiner. For Earman and D. Fraser, the focus is on the existence of a unitary intertwiner between the free field and the unrenormalized interacting field. In this case, an IR cutoff suffices to defuse Haag’s theorem and ensure the existence of an unitary intertwiner. For Klaczynski and Maiezza and Vasquez, the focus is on the existence of a unitary intertwiner between the free field and the renormalized interacting field. In this case, substantive engagement with renormalization techniques will be necessary for settling the question of whether a satisfactory intertwiner exists. Thus, applying our framework to this debate over the IP reveals that the real debate to be had is over the exact version of the fields for which we are in want of an intertwiner.
The framework therefore reveals two substantial choices underlying the superficial disagreement about the IP’s existence. First, must we explicitly identify a non-perturbative structure in order to be satisfied with our theory’s representation of scattering phenomena? Second, why, hence where, do we want a well-defined unitary intertwiner? In particular, is it enough for an intertwiner to relate the (IR safe and scale-relative) observables remaining after renormalization to appropriate free-field correlates (i.e., unitary in that amplitudes appropriately coincide), or do we want an intertwiner to find correlates for the entire free-field state space and observable structure in a representation of the interacting field (i.e., unitary in that it is a dynamical operator defined on a single state space)?
@@
The reader might think Kastner fits in with Bain. This is not correct, however. Kastner not only recommends a different interpretation of particle physics from Halvorson and Clifton but, as her proposed renovations reveal, suggests QFT must be replaced altogether. That is, rather than derive an interpretation from a presupposed formalism, she instead derives a formalism from a presupposed interpretation. Among other things, this interpretation is committed to a particle conception of particle physics. This commitment ultimately derives from a belief that relativistic quantum mechanics should behave more non-locally than QFT does, as, she thinks, non-relativistic quantum mechanics counsels. This suggests two disanalogies with Bain. First, the comparatively unmoored approximations of canonical QFT are not enough—we should demand exact calculations, or at least approximations to exact solutions. Second, a fully satisfactory theory should wear its metaphysics on its face more plainly than does canonical QFT.
As above, there are substantial choices underlying the superficial disagreement about particle existence. First, is the conception of particles inherited from the free-field formalism a satisfactory one for interacting QFT? Second, at what point should we be satisfied with a mathematical formalism in its representational capacity? In particular, is it (at least potentially) metaphysically significant that the calculations delivered by canonical QFT are not approximations to exact solutions?
@@
Spelling out the renovations proposed also reveals commitments we might have otherwise missed (see figure 2). Focusing just on broad strokes aims as a way to separate various responses to Haag’s theorem, Kastner and Seidewitz would seem far removed from the puzzle-focused theory interpreters—the former propose new theoretical forms for QFT, while these interpreters restrict themselves to the formal variants. However, it is worth noting that these groups of authors seem to share the same outlook on axiomatization. In particular, both seem at least weakly committed to claiming that QFT phenomena has a structure that can be captured in a closed form as something like a finite set of axioms and that we can actually specify these axioms. Likewise, the puzzle-focused interpreters who believe particles have been thoroughly banished believe not only that QFT can be so captured but that it has. In contrast, Earman and D. Fraser appear committed only to the belief that axiomatization could be helpful for comprehending current QFT, or even for finding a more scrutable form. D. Fraser, at least, is quite clear that CQFT does not need to build a realistic model in order to have been useful—its utility lies rather in the fleshing-out of the relationship among the axioms. While Kastner and Seidewitz certainly could take such a view, the emphasis they place on the correctness of their chosen axioms suggests this is not their current conception.
@@
This is useful, too, for understanding the varied responses to the question of which assumption of Haag’s theorem should be dropped. Kastner, Seidewitz, and the most staid puzzle-focused interpreters each have a clear answer to this question. Things get murkier as we consider the others, however. Not only do none of them explicitly affirm some set of precise assumptions, they are not entirely clear about which assumptions are problematic and how. This comes in degrees, of course. Earman and D. Fraser are perhaps clearest, suggesting that the IP’s assumption of a global unitary intertwiner is problematic but that the local equivalence delivered by Haag-Ruelle appears to suffice. Klaczynski and Maiezza and Vasquez are a bit less clear. Insofar as they each hope for an explicit formalism, unlike canonical QFT, they appear sympathetic to axiomatic approaches. Klaczynski only suggests that renormalization somehow replaces the unitarity of the intertwiner.
Maiezza and Vasquez are agnostic regarding the question of which assumptions leading to Haag’s theorem should be dropped. In their proposed model DRQFT, they drop the assumption of a Poincaré-invariant vacuum state, but they present this as only one possible and promising path forwards. [^45] Thus they propose relaxing the same assumption as Seidewitz. However, their model is expressed predominantly in the language of canonical, non-axiomatic QFT. This model is intended to encode unknown physics at high energy scales. [Maiezza and Vasquez, personal communication, Feb. 2, 2024]
In a sense, Duncan is even less clear, albeit for a principled reason. Obviously respectful of axiomatization’s utility, he nevertheless emphasizes again and again that the truly remarkable feature of QFT is the “scale separation” property of local field theory [Duncan, 2012, 570-1]. Building one’s theory around this fact, as the effective Lagrangian approach of canonical QFT has done, would seem to undermine specification of any single set of assumptions since, e.g., symmetry structure can vary. Duncan, it would seem, cannot be clearer about what is denied.
Here our framework, by distinguishing Assessments and Repairs from Renovations and Extra-Haagian Outlooks, calls our attention to a whole host of choices underlying the superficial disagreement. In general, these concern the utility and feasibility of formally-driven approaches to expanding QFT. In particular, what is the point of axiomatization? what can we reasonably expect it to afford us? how does (should) it interface with other programs in QFT?
@@ 5.2 Lessons for making progress
In the first instance, the framework we have provided constitutes guidance for highlighting the choices driving the more superficial disagreements. But this is not its most important guidance. More dramatically, we think the framework highlights a kind of question that philosophers of physics often overlook regarding the so-called foundations.
Many are tempted to ask, as we seem to do in the title: how foundational is Haag’s theorem to QFT, really? The adverb signals a temptation to think there is a universal extra-Haagian outlook on QFT. Of course, as we see it, there is no such outlook. Yes, there is wide agreement that QFT long ago secured much comfortable space to wander around. But we have not rested there—we wish to expand these rooms, and in so doing we soon face choices about how best to secure our foundations. Many putative problems have arisen from attempts to do so, Haag’s theorem among them. Is evading Haag’s theorem central to all of these expansions? That is, does Haag’s theorem need to be explicitly centered in all expansion programs—say, by setting its evasion as a derivational goal, or by cementing involved concepts as essential to the representation of QFT? Surely not—even Hilbert recognized that such designations were relative to the proposed expansion and, at any rate, only tentative [Hilbert, 1918]. If an expansion does not call for, say, securing global unitary equivalence between free and interacting fields, or for Poincaré covariance of the full state space, or unambiguous analytic continuation to finite coupling, or some other such security, Haag’s theorem simply is not relevant.
Nevertheless, there are expansions for which Haag’s theorem is relevant, so one can easily ask the same question in a localized form: how foundational is Haag’s theorem for this expansion? This seems a more sensible question, and in fact it seems to be asking several sensible questions. Most straightforwardly, it could be asking how radically the form of QFT needs to change in order to secure the proposed expansion against the ravages of Haag’s theorem. The radicalness of Kastner’s and Seidewitz’s expansions stand out here, suggesting that Haag’s theorem is profoundly foundational for their expansions; on the opposite end, Duncan and Miller propose very little in the way of expansion, and Haag’s theorem therefore does not appear foundational since no change is necessary. Call questions like this formal questions about how foundational Haag’s theorem is. This seems to us to be how most philosophers of physics would approach grading how foundational a (cluster of) results is: the more ‘central’ a concept or statement is within the theoretical apparatus in some formal sense—cashed out either semantically or syntactically—the more foundational it is.
Yet it is more difficult to answer the formal question about how foundational Haag’s theorem is with respect to the other proposed expansions. This is because these expansion projects are actively probing the theoretical and mathematical possibilities for achieving their goals and, as such, Haag’s theorem does not have as definitive a formal status. But this reveals a set of methodological questions you might have in mind when asking about how foundational Haag’s theorem is. Each of the expansion projects takes a position on the role of reasoning done with more abstract or idealized assumptions, more exacting characterizations of representations, or, in general, at farther remove from any particular, concrete modeling situation. Call this foundational reasoning, to distinguish it from the (putatively foundational) concepts or statements of the last paragraph. Note that the two senses of foundational need not coincide: ultimately Haag’s theorem may not require much adjustment to address (formal sense), but clearly the reasoning that led to it, and that expanded around it, is foundational in this methodological sense.
Rather than inquire about the formal status of Haag’s theorem, we could inquire about reasoning of its ilk. How fruitful was the reasoning that led to Haag’s theorem? Is such reasoning still useful, so that it remains deserving of attention? Why or why not? Such questions are methodological insofar as they primarily concern the nature of the growth of human knowledge rather than the nature of its currently supposed contents. In more general terms, we might ask:
@@
    What role does (should) foundational reasoning play in progress in physics?
@@
    How does (should) foundational reasoning coordinate with non-foundational work?
@@
    And, what does (should) foundational reasoning even look like?
Asking these methodological questions opens up a comparatively new avenue of inquiry for philosophers of physics by shifting our attention from the past—what is the formal-foundational status of Haag’s theorem—to the future—how fruitful will Haag’s theorem-related reasoning be.
And here is the most significant contribution of the framework: in attempting to decide on the status of our QFT knowledge vis-á-vis Haag’s theorem, we have been dragged into a decision context where we must forecast the future of the various proposals. In Nickle’s terms [Nickles, 2006], we have been dragged from philosophers’ usual concern for epistemic appraisal—is it well-confirmed or truth-conducive?—to a concern for heuristic appraisal—does it have a good “opportunity profile”? Our positions is that, when it comes to thinking hard about our best physical theories, being dragged in this way is neither new nor unwelcome. Indeed, as Streater and Wightman humorously (but correctly) observe, many physicists felt that the development of a more rigorous quantum field theory was supremely ill-advised; cynics had even “compared them to the Shakers, a religious sect of New England who built solid barns and led celibate lives, a non-scientific equivalent of proving rigorous theorems and calculating no cross sections” [Streater and Wightman, 2000, 1]. That is, the “opportunity profile” of (rigorous) foundational reasoning about quantum field theory was considered weak.
The early negative assessment of rigorous approaches to QFT was born of beliefs about the comparatively greater promise of investigating the classical foundations of field theory versus the foundations of a quantum field theory. And yet, many of the results that give comfort in the face of Haag’s theorem—Reed and Simon’s Theorem X.46 and the contemporary understanding of LSZ among them—were at least informed by the quantum Shakers above who aimed to “kill it or cure it”, as Streater and Wightman put it [Streater and Wightman, 2000]. So, does this vindicate the kind of foundational reasoning Shaker-physicists, such as Haag or Wightman, engaged in? And if it does, does it tell us anything about the auspiciousness of such reasoning today? These are the kinds of methodological questions to which our attention has been drawn by the framework’s centering of scientific reasoning rather than individual concepts or claims.
We think asking methodological questions about foundational reasoning better equips philosophers for engaging with ongoing inquiry. After all, our respect of science does not boil down only to our trust of the propositional claims scientists make about what the world is like. Rather, we presumably trust these claims because we trust science is composed of experts “knowing how to conduct research, including HA [heuristic assessment] of the various proposals for fruitful inquiry” [Nickles, 2015, 84]. In particular, we trust that they have the expertise necessary to decide which theorems or concepts it would be fruitful to place at the foundation of a theory.
[…]
Nevertheless, serendipity is not an especially defensible plan for pursuing future physics, which leaves us to decide among the kinds of foundational reasoning currently on offer. Supposing we have set ourselves the goal of specifying an unambiguous structure underlying canonical QFT’s perturbative calculations, how should we go about doing so? As the pluralist response we began this section with emphasized, it seems likely that a mixed, rather than pure, strategy will be appropriate. We simply do not know definitively which program(s) will bear fruit. But the question is, which mixed strategy? This is where it would be helpful to compare and contrast the extant programs, considering especially how they complement or compete against alternatives as well as the reasons given for or against their auspiciousness. Should these reasons be historical or abstracted away from the particulars of the problem of, say, Haag’s theorem, there is all the more reason for philosophers of science to get involved.
It may seem that we are calling for a retread of the path worn by the so-called D. Fraser-Wallace debate [Wallace, 2006, Wallace, 2011, Fraser, 2011, Fraser, 2009] (see also [Koberinski, 2016, Koberinski, 2022] for reception of the debate). There is something to this charge. Both D. Fraser and Wallace have views on the auspiciousness of various proposed expansions of QFT, and later work from both comes to consider questions like those we pose above more directly (e.g., [Fraser, 2022, Wallace, 2020]). However, the debate as it exists in print is ultimately concerned with what form of QFT is worth subjecting to philosophical methods of interpretation, as well as what those methods should be. It is not (explicitly) about the nature or role of foundational reasoning for physics, let alone about the various goals physicists and mathematicians have set for themselves, the auspiciousness of the programs they champion, or the reasons they give for such assessments of auspiciousness.
The present work is better recognized as calling for extending a path concerned with the scientific value of different kinds of foundational reasoning. [Koberinski, 2023, 14], for example, argues that the axiomatic-adjacent approach of framework generalizations is reasonable insofar as it “expand[s] the domain of indirect tests able to find evidence of new physics.” In another vein, [Mitsch, 2022] argues that no-go theorems can anchor mathematical research programs in serious interpretive issues. [Freeborn et al., 2023] argue, moreover, that no-go theorems can serve as helpful guides to building adequate models. We are suggesting that further work on the utility of the various conceptual apparatuses, methods, exploratory techniques, etc., on offer could benefit scientific reasoning about what projects to take up and, thereby, the choices made. Figure 2 provides a loose disciplinary sort of some of the programs we encountered in our discussion of Haag’s theorem; each is arguably engaged foundational reasoning, which makes them good targets for further investigation along this path. In doing so, we should keep two things in mind. First, QFT is not a methodological island: there is a rich history of physics—and indeed of other sciences!—on which we can draw as we adumbrate the virtues and vices of the various reasoning tools in our tool belt. And second, we must be ready to adapt when we inevitably face new goals, pragmatic constraints, or tool adaptations, and the like.